OpenStudy (anonymous):

6^3m times 6^-m=6^-2m

OpenStudy (anonymous):

Do you have choices I might be able to narrow it down for you

OpenStudy (anonymous):

No, I have to solve it. I know the answer is 0, I just don't know how my teacher got 0

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

\(\large \bf { 6^{3m}\cdot 6^{-m}=6^{-2m}\implies {\color{brown}{ 6}}^{3m+(-m)}={\color{brown}{ 6}}^{-2m}\qquad meaning \\ \quad \\ 3m+(-m)=-2m }\)

OpenStudy (jdoe0001):

same base, for the EQUATion to be true the exponents MUST also be equal

OpenStudy (anonymous):

How do I make them equal to each other?